OpenStudy (anonymous):

A sail is in the form of a right triangle that is three times as high as its wide. The sail is made from 6 square meters of material. What is its height?

OpenStudy (anonymous):

width=2 and height=6

OpenStudy (anonymous):

Area of a Triangle = 1/2 * bh The base can be x and the height 3x So, 6 = 1/2 * x(3x) 6 = 1/2 * 3x^2 12 = 3x^2 4 = x^2 x = 2